# Improving Generalization of Neural Combinatorial Optimization for Vehicle Routing Problems via Test-Time Projection Learning

### v2

@@ 1 Introduction
The vehicle routing problem (VRP) is prevalent in critical domains such as transportation [1], logistics [2], and supply chain management [3]. However, due to the NP-hard nature, solving a VRP can be complex and time-consuming [4]. While traditional exact and heuristic methods can yield optimal or near-optimal solutions, their high computational cost and dependence on domain expertise consequently diminish their feasibility for real-world deployment.
@@
In recent years, neural combinatorial optimization (NCO) has emerged as a promising approach for solving VRP [5]. It utilizes neural networks to learn problem-solving strategies directly from the VRP instances, minimizing the need for algorithms designed by experts. Once trained, these neural networks can efficiently generate solutions for new instances at a low computational cost. Therefore, NCO methods demonstrate potential advantages in tackling limitations of traditional methods and exhibit promising performance in solving small-scale VRPs [6, 7, 8, 9].
@@
However, due to the different distributions between small and large-scale instances, the effectiveness of existing approaches degrades substantially when it comes to large-scale problems (e.g., problems with more than 1K nodes), thereby severely limiting their practical capability. To address the scalability issue, some attempts have been devoted to training NCO models on larger VRP instances (i.e., instances with 500 nodes) [10, 11]. However, the existing supervised learning (SL) and reinforcement learning (RL) both show their shortcomings when training on large-scale VRP instances. SL lacks sufficient labels (e.g., high-quality solutions) while RL suffers from extremely sparse rewards.
@@
Consequently, current methods focus on training a model on small-scale instances and generalizing to large-scale scenarios via decomposition or local policy [12, 13, 14, 15]. The decomposition-based method first simplifies large-scale problems to a series of subproblems. Subsequently, a solver trained on the small-scale instances can be utilized to construct a partial solution of the subproblems [16, 17, 18]. However, the scope of the decomposition often requires manual tuning, and even decomposing the problem could change its property, and thus damage the optimality of the solving algorithm. Another effort is local policy-based methods [14, 15]. It first reduces the search space to a small candidate subgraph at each construction step, typically based on the Euclidean distance to the last visited node. The next node is decided by the original policy or a local policy.
@@
However, the distribution of selected subgraphs often differs from training instances, which significantly impairs the model’s scalability. Current methods employ a projection technique to effectively transform these varied input distributions into a uniform distribution encountered during training [15, 13, 19]. Nevertheless, existing strategies require integration with the model training to ensure effectiveness during testing. Meanwhile, these manually designed strategies heavily rely on specialized domain expertise, thereby limiting their adaptability.
@@ 2 Related Work
Existing NCO methods can be categorized into constructive [20, 21, 22, 23, 24, 25] and non-constructive [26, 27, 28, 29]. This paper mainly focuses on the constructive method; a detailed review of the non-constructive method and other LLM-based methods can be found in the Appendix. A.
@@ 2.1 Direct Generalization
This kind of method often trains neural networks on small-scale instances and generalizes them to large-scale instances. It is mostly known as the construction-based methods that predict an approximate solution in an autoregressive manner. Early studies demonstrate that neural models trained with SL or RL can reach promising performance on small-scale instances [30, 31, 32]. Moreover, [6, 33] propose to use the Transformer architecture [34] to design powerful attention-based models to solve VRPs. Subsequently, various Transformer-based approaches have been developed with different strengths [35, 36, 37, 38, 39, 40, 41]. Meanwhile, many studies focus on improving model performance on large-scale VRPs [42]. In which [42] and [18] train the model with SL on 100-node and exhibit generalization ability to 1K nodes. Specifically, BQ [42] modifies the Markov Decision Process (MDP) to efficiently leverage the symmetries of combinatorial optimization problems (COPs). LEHD [18] develops a light encoder and heavy decoder structure to reach the same result. Nevertheless, the training distribution of 100-node instances drastically differs from the large-scale instances; the trained model usually performs poorly when tested on this large-scale scenario. Consequently, several attempts have trained models on larger-scale instances with up to 500 nodes [43, 11, 38, 10] or even directly training on the large-scale [44] (up to 100K-nodes). However, these approaches introduce prohibitive computational costs due to the exponentially growing search space.
@@ 2.2 Decomposition-Based Generalization
Apart from directly generalizing the model trained on a small-scale dataset to the large-scale instances, another line of research focuses on decomposing the large node graph into several small-scale subproblems. This approach involves solving them separately with a particular solver and then combining their solutions to form the solution of the large-scale problem [45, 46, 47, 12, 13, 48]. In particular, the problem decomposition and subproblem solving are often involved in different models. The task of the subproblem solver can be constructing a complete solution of a small-scale VRP or partial solutions of large-scale instances [16, 17, 18]. However, such decomposition introduces two critical shortcomings. When solving other complicated VRPs (e.g., CVRP), such decomposition becomes intractable and cannot be achieved by an individual strategy or model. Furthermore, when the decomposed subproblem is not small enough, it may still require other generalization techniques to solve.
@@ 2.3 Local Policy-Based Generalization
Another approach to solving large-scale VRPs is to reduce the search space to the $K$-Nearest-Neighbors (KNN) from the last visited node. The final decision is guided by either the original neural model with auxiliary distance information [49] or a local policy [14]. In addition, [15, 50, 42] directly select the candidate node from the neighborhood using NCO models. Although local policy can efficiently reduce the search space for constructive NCO, the reduced subgraph often drastically differs from the training instances, especially in large-scale or non-uniform instances. Consequently, the NCO models fail to distinguish the promising node. To this end, [15, 13, 19] propose a projection strategy to transform the extracted subgraph to a uniform distribution. While these strategies can enhance generalization, their effectiveness typically requires integration into model training, thereby imposing additional computational overhead and limiting their adaptability.
@@ 3 Preliminaries
where $f_{T}(x,h)$ quantifies the effectiveness of heuristic $h$ on instance $x$. The heuristic space $H$ encompasses all feasible strategies that adhere to the constraints of $T$. This approach aligns with hyper-heuristic methodologies [51], which automate heuristic selection or generation without domain-specific handcrafting.
@@
LLM-based AHD integrates LLMs into the evolutionary search for a high-performance heuristic [52, 53, 54, 55]. This paradigm leverages LLMs to generate and refine heuristics through natural language reasoning and code synthesis. Within an evolutionary computation framework, LLMs simulate mutation and crossover operations by modifying or combining existing heuristics. For example, given a parent heuristic, LLMs can introduce new algorithmic components, adjust parameters, or merge features from multiple candidates, guided by linguistic prompts that encode task-specific objectives and constraints. This iterative process balances the exploration of high-quality candidates, enabling efficient traversal of the heuristic space $H$.
@@ 4 Distribution Projection in Large-Scale VRPs
Figure 1: (a) Solution construction process for solving TSP5K instance, gray node, blue node, red node, and grey line denote the unvisited nodes, KNN (kk=100) of the current node, current node, and constructed partial route, respectively. (b) Extracted KNN graph, RWD [56] indicates the distance between the input graph and a uniformly distributed training instance. (c) Projected KNN graph. (d) Training instance with 100 nodes.
When solving large-scale VRPs, current methods are facing challenges from the exponentially growing search space. As shown in the Figure. 1 (a), existing works reduce the search space to the KNN of the last visited node. Despite this reduction providing sufficient improvement, Figure. 1 (b) demonstrates that the selected KNN subgraphs come with various distributions that drastically differ from the training instance. Making it difficult for the model to predict the next promising node. Recent studies provide several projection techniques to project the input subgraph to a uniform distribution [15, 13, 19]. In this section, we systematically analyze the effectiveness of these projection methods and observe that such a technique has a great impact on the model’s final performance.
@@
To validate the effectiveness of the projection, we employ the LEHD [18] as the base model, restricting its search space to the KNN ($k=100$) nodes from the last visited node. Then, we apply the projection method used in INViT [15] to map the subgraph. Subsequently, the LEHD is adopted to select one promising node within the transformed subgraph. The experimental results are provided in Table 1. The experiment results exhibit that LEHD with the INViT projection method outperforms LEHD without projection consistently on all scales in TSP.
@@ 5.1 Test-Time Projection Learning
14:   Generate new offspring using prompt strategies from [52]
@@
The overall framework of TTPL is demonstrated in the Algorithm. 1. Its inputs consist of the population size $N$, number of generations $N_{g}$, the NCO model $f_{\theta}$, the LLM used to evolve the strategy, and the evaluation data set $D$. TTPL follows the general framework of EoH [52]. Therefore, we first generate $N$ individuals using LLM with the prompts from EoH. Here, individuals consist of a description in natural language and a code block that implements the projection. These individuals are then evaluated by the NCO model $f_{\theta}$ to construct the initial population $P_{0}$. Moreover, we conduct $N_{g}$ iterations to obtain an optimal projection strategy. In each iteration, we prompt LLMs to design different individuals based on a set of selected parents. Once the new population is generated, we select the first $N$ best individuals as the next generation. Finally, it outputs the optimal strategy it found in previous iterations
@@
where $r_{i}$ is the rank of their fitness. After that, the selected parents are used to generate offspring through the four prompt strategies in [52], which consist of two groups, namely, Exploration (E1, E2) and Modification (M1, M2). The exploration prompt focuses on generating new individuals based on several selected parent strategies, while the modification prompt aims to modify the selected strategy for better performance. The details of four prompt strategies can be found in the Appendix. B.1
@@ 5.2 Multi-View Decision Fusion
In particular, given a subgraph with coordinates $\mathbf{X}\in\mathbb{R}^{k\times 2}$, we generate $M$ augmented variants through geometric transformations $\mathcal{T}_{m}(\mathbf{X})$ proposed in [6]. Each augmented subgraph $\mathbf{X}_{m}=\mathcal{T}_{m}(\mathbf{X})$ is solved by the NCO model to produce logits $\mathbf{l}_{m}\in\mathbb{R}^{K}$. The final selection probability is derived by aggregating logits across all augmentations:
@@ 6 Experiment
We evaluate our method on TSP and CVRP instances. Following established benchmarks [47, 18], the test sets include 128 instances for TSP1K and 16 instances each for TSP5K, 10K, 50K, and 100K. For CVRP, we use 100 instances for CVRP1K and CVRP5K and 16 instances each for CVRP10K, CVRP50K, and CVRP100K. CVRP capacities are 200 for CVRP1K and 300 for CVRP5K/10K/50K/100K. Optimal TSP solutions are calculated using LKH3 [57], while CVRP solutions are derived using HGS [58]. To further assess generalization, we follow [44] and evaluate on real-world instances from TSPLIB [59] and CVRPLIB [60], selecting symmetric EUC_2D instances with over 1K nodes (33 TSP and 14 CVRP instances). This ensures alignment with both synthetic benchmarks and practical scenarios.
@@
Our experiments adopt LEHD as the base model [18], aligning with its original architecture and hyperparameters. The embedding dimension is configured to 128, and the decoder comprises 6 attention layers. Each multi-head attention layer uses 8 heads, while the feed-forward layer dimension is set to 512. We configure the KNN to 100. All experiments are performed on a single NVIDIA GeForce RTX 3090 GPU with 24GB of memory.
@@
The evolutionary framework follows the approach proposed by EoH [52]. The population size is set to 20, and the evolution process is run for 105 generations. The parent heuristic is set to 2. The GPT-4o-mini model serves as the LLM backbone. For TSP/CVRP 1K, 5K, and 10K, we optimize one projection strategy for each scale, and generalize the strategy optimized on 10K to TSP/CVRP 50K and 100K due to the computational burden. At the beginning of the optimization, we input the projection strategy proposed in [15] as a seed to guide the heuristic design.
@@
We compare TTPL with six categories of methods: Classical Solvers: Concorde [61], LKH3 [57] and HGS [58]; Insertion Heuristic: Random Insertion; Construction-based NCO Methods: POMO [7], BQ [42], LEHD [18], INVIT [15] and SIGD [62]; Heatmap-based Methods: DIFUSCO [28]; Decomposition-based Method: GLOP [13], H-TSP [12]; Local Construction-based Method: ELG [14].
@@
We evaluate the performance using the optimality gap and inference time. The optimality gap quantifies the percentage difference between solutions generated by our method and the ground truth obtained from LKH3 (TSP) or HGS (CVRP). Inference time, reported in seconds (s), minutes (m), and hours (h), captures the computational efficiency of generating solutions across varying problem scales. For TTPL, we report three types of results, those obtained by greedy trajectory with a purely designed projection method, those obtained by greedy trajectory with our designed projection and MVDF, and those obtained with projection, MVDF, and Random Re-Construction (RRC) under 1000 iterations [18].

### v3

@@ 1 Introduction
The vehicle routing problem (VRP) is prevalent in critical domains such as transportation [veres2019deep], logistics [laterre2018ranked], and supply chain management [giallanza2020fuzzy]. However, due to the NP-hard nature, solving a VRP can be complex and time-consuming [ausiello2012complexity]. While traditional exact and heuristic methods can yield optimal or near-optimal solutions, their high computational cost and dependence on domain expertise consequently diminish their feasibility for real-world deployment.
@@
In recent years, neural combinatorial optimization (NCO) has emerged as a promising approach for solving VRP [bengio2021machine]. It utilizes neural networks to learn problem-solving strategies directly from the VRP instances, minimizing the need for algorithms designed by experts. Once trained, these neural networks can efficiently generate solutions for new instances at a low computational cost. Therefore, NCO methods demonstrate potential advantages in tackling limitations of traditional methods and exhibit promising performance in solving small-scale VRPs [kool2018attention, kwon2020pomo, hottung2021efficient, zhou2025urs].
@@
However, due to the different distributions between small and large-scale instances, the effectiveness of existing approaches degrades substantially when it comes to large-scale problems (e.g., problems with more than 1K nodes), thereby severely limiting their practical capability. To address the scalability issue, some attempts have been devoted to training NCO models on larger VRP instances (i.e., instances with 500 nodes) [jin2023pointerformer, zhou2023towards]. However, the existing supervised learning (SL) and reinforcement learning (RL) both show their shortcomings when training on large-scale VRP instances. SL lacks sufficient labels (e.g., high-quality solutions) while RL suffers from extremely sparse rewards.
@@
Consequently, current methods focus on training a model on small-scale instances and generalizing to large-scale scenarios via decomposition or local policy [pan2023h, ye2024glop, gao2023towards, fang2024invit]. The decomposition-based method first simplifies large-scale problems to a series of subproblems. Subsequently, a solver trained on the small-scale instances can be utilized to construct a partial solution of the subproblems [kim2021learning, cheng2023select, luo2023neural]. However, the scope of the decomposition often requires manual tuning, and even decomposing the problem could change its property, and thus damage the optimality of the solving algorithm. Another effort is local policy-based methods [gao2023towards, fang2024invit]. It first reduces the search space to a small candidate subgraph at each construction step, typically based on the Euclidean distance to the last visited node. The next node is decided by the original policy or a local policy.
@@
However, the distribution of selected subgraphs often differs from training instances, which significantly impairs the model’s scalability. Current methods employ a projection technique to effectively transform these varied input distributions into a uniform distribution encountered during training [fang2024invit, ye2024glop, fu2021generalize]. Nevertheless, existing strategies require integration with the model training to ensure effectiveness during testing. Meanwhile, these manually designed strategies heavily rely on specialized domain expertise, thereby limiting their adaptability.
@@ 2 Related Work
Existing NCO methods can be categorized into constructive [huang2025rethinking, kwon2021matrix, drakulic2024goal, liu2024multi, zhou2024mvmoe, khalil2017learning] and non-constructive [zhang2025adversarial, li2023t2t, sun2023difusco, li2024fast]. This paper mainly focuses on the constructive method; a detailed review of the non-constructive method and other LLM-based methods can be found in the Appendix. A.
@@ 2.1 Direct Generalization
This kind of method often trains neural networks on small-scale instances and generalizes them to large-scale instances. It is mostly known as the construction-based methods that predict an approximate solution in an autoregressive manner. Early studies demonstrate that neural models trained with SL or RL can reach promising performance on small-scale instances [vinyals2015pointer, bello2016neural, nazari2018reinforcement]. Moreover, [kool2018attention, deudon2018learning] propose to use the Transformer architecture [vaswani2017attention] to design powerful attention-based models to solve VRPs. Subsequently, various Transformer-based approaches have been developed with different strengths [xin2020step, sun2024learning, zheng2024dpn, zhou2024instance, zheng2024Pareto_Improver, li2025cada, luo2025rethink_tightness]. Meanwhile, many studies focus on improving model performance on large-scale VRPs [drakulic2023bq]. In which [drakulic2023bq] and [luo2023neural] train the model with SL on 100-node and exhibit generalization ability to 1K nodes. Specifically, BQ [drakulic2023bq] modifies the Markov Decision Process (MDP) to efficiently leverage the symmetries of combinatorial optimization problems (COPs). LEHD [luo2023neural] develops a light encoder and heavy decoder structure to reach the same result. Nevertheless, the training distribution of 100-node instances drastically differs from the large-scale instances; the trained model usually performs poorly when tested on this large-scale scenario. Consequently, several attempts have trained models on larger-scale instances with up to 500 nodes [cao2022dan, zhou2023towards, zhou2024instance, jin2023pointerformer] or even directly training on the large-scale [luoboosting] (up to 100K-nodes). However, these approaches introduce prohibitive computational costs due to the exponentially growing search space.
@@ 2.2 Decomposition-Based Generalization
Apart from directly generalizing the model trained on a small-scale dataset to the large-scale instances, another line of research focuses on decomposing the large node graph into several small-scale subproblems. This approach involves solving them separately with a particular solver and then combining their solutions to form the solution of the large-scale problem [li2021learning, zong2022rbg, hou2023generalize, pan2023h, ye2024glop, zheng2024udc]. In particular, the problem decomposition and subproblem solving are often involved in different models. The task of the subproblem solver can be constructing a complete solution of a small-scale VRP or partial solutions of large-scale instances [kim2021learning, cheng2023select, luo2023neural]. However, such decomposition introduces two critical shortcomings. When solving other complicated VRPs (e.g., CVRP), such decomposition becomes intractable and cannot be achieved by an individual strategy or model. Furthermore, when the decomposed subproblem is not small enough, it may still require other generalization techniques to solve.
@@ 2.3 Local Policy-Based Generalization
Another approach to solving large-scale VRPs is to reduce the search space to the $K$-Nearest-Neighbors (KNN) from the last visited node. The final decision is guided by either the original neural model with auxiliary distance information [wang2024distance] or a local policy [gao2023towards]. In addition, [fang2024invit, zhou2025l2r, drakulic2023bq] directly select the candidate node from the neighborhood using NCO models. Although local policy can efficiently reduce the search space for constructive NCO, the reduced subgraph often drastically differs from the training instances, especially in large-scale or non-uniform instances. Consequently, the NCO models fail to distinguish the promising node. To this end, [fang2024invit, ye2024glop, fu2021generalize] propose a projection strategy to transform the extracted subgraph to a uniform distribution. While these strategies can enhance generalization, their effectiveness typically requires integration into model training, thereby imposing additional computational overhead and limiting their adaptability.
@@ 3 Preliminaries
where $f_{T}(x,h)$ quantifies the effectiveness of heuristic $h$ on instance $x$. The heuristic space $H$ encompasses all feasible strategies that adhere to the constraints of $T$. This approach aligns with hyper-heuristic methodologies [burke2013hyper], which automate heuristic selection or generation without domain-specific handcrafting.
@@
LLM-based AHD integrates LLMs into the evolutionary search for a high-performance heuristic [liu2024evolution, zheng2025monte, li2025ars, li2025multi]. This paradigm leverages LLMs to generate and refine heuristics through natural language reasoning and code synthesis. Within an evolutionary computation framework, LLMs simulate mutation and crossover operations by modifying or combining existing heuristics. For example, given a parent heuristic, LLMs can introduce new algorithmic components, adjust parameters, or merge features from multiple candidates, guided by linguistic prompts that encode task-specific objectives and constraints. This iterative process balances the exploration of high-quality candidates, enabling efficient traversal of the heuristic space $H$.
@@ 4 Distribution Projection in Large-Scale VRPs
Figure 1: (a) Solution construction process for solving TSP5K instance, gray node, blue node, red node, and grey line denote the unvisited nodes, KNN (kk=100) of the current node, current node, and constructed partial route, respectively. (b) Extracted KNN graph, RWD [wangefficient] indicates the distance between the input graph and a uniformly distributed training instance. (c) Projected KNN graph. (d) Training instance with 100 nodes.
When solving large-scale VRPs, current methods are facing challenges from the exponentially growing search space. As shown in the Figure. 1 (a), existing works reduce the search space to the KNN of the last visited node. Despite this reduction providing sufficient improvement, Figure. 1 (b) demonstrates that the selected KNN subgraphs come with various distributions that drastically differ from the training instance. Making it difficult for the model to predict the next promising node. Recent studies provide several projection techniques to project the input subgraph to a uniform distribution [fang2024invit, ye2024glop, fu2021generalize]. In this section, we systematically analyze the effectiveness of these projection methods and observe that such a technique has a great impact on the model’s final performance.
@@
To validate the effectiveness of the projection, we employ the LEHD [luo2023neural] as the base model, restricting its search space to the KNN ($k=100$) nodes from the last visited node. Then, we apply the projection method used in INViT [fang2024invit] to map the subgraph. Subsequently, the LEHD is adopted to select one promising node within the transformed subgraph. The experimental results are provided in Table 1. The experiment results exhibit that LEHD with the INViT projection method outperforms LEHD without projection consistently on all scales in TSP.
@@ 5.1 Test-Time Projection Learning
14:   Generate new offspring using prompt strategies from [liu2024evolution]
@@
The overall framework of TTPL is demonstrated in the Algorithm. 1. Its inputs consist of the population size $N$, number of generations $N_{g}$, the NCO model $f_{\theta}$, the LLM used to evolve the strategy, and the evaluation data set $D$. TTPL follows the general framework of EoH [liu2024evolution]. Therefore, we first generate $N$ individuals using LLM with the prompts from EoH. Here, individuals consist of a description in natural language and a code block that implements the projection. These individuals are then evaluated by the NCO model $f_{\theta}$ to construct the initial population $P_{0}$. Moreover, we conduct $N_{g}$ iterations to obtain an optimal projection strategy. In each iteration, we prompt LLMs to design different individuals based on a set of selected parents. Once the new population is generated, we select the first $N$ best individuals as the next generation. Finally, it outputs the optimal strategy it found in previous iterations
@@
where $r_{i}$ is the rank of their fitness. After that, the selected parents are used to generate offspring through the four prompt strategies in [liu2024evolution], which consist of two groups, namely, Exploration (E1, E2) and Modification (M1, M2). The exploration prompt focuses on generating new individuals based on several selected parent strategies, while the modification prompt aims to modify the selected strategy for better performance. The details of four prompt strategies can be found in the Appendix. B.1
@@ 5.2 Multi-View Decision Fusion
In particular, given a subgraph with coordinates $\mathbf{X}\in\mathbb{R}^{k\times 2}$, we generate $M$ augmented variants through geometric transformations $\mathcal{T}_{m}(\mathbf{X})$ proposed in [kool2018attention]. Each augmented subgraph $\mathbf{X}_{m}=\mathcal{T}_{m}(\mathbf{X})$ is solved by the NCO model to produce logits $\mathbf{l}_{m}\in\mathbb{R}^{K}$. The final selection probability is derived by aggregating logits across all augmentations:
@@ 6 Experiment
We evaluate our method on TSP and CVRP instances. Following established benchmarks [hou2023generalize, luo2023neural], the test sets include 128 instances for TSP1K and 16 instances each for TSP5K, 10K, 50K, and 100K. For CVRP, we use 100 instances for CVRP1K and CVRP5K and 16 instances each for CVRP10K, CVRP50K, and CVRP100K. CVRP capacities are 200 for CVRP1K and 300 for CVRP5K/10K/50K/100K. Optimal TSP solutions are calculated using LKH3 [helsgaun2017extension], while CVRP solutions are derived using HGS [vidal2022hybrid]. To further assess generalization, we follow [luoboosting] and evaluate on real-world instances from TSPLIB [reinelt1991tsplib] and CVRPLIB [uchoa2017new], selecting symmetric EUC_2D instances with over 1K nodes (33 TSP and 14 CVRP instances). This ensures alignment with both synthetic benchmarks and practical scenarios.
@@
Our experiments adopt LEHD as the base model [luo2023neural], aligning with its original architecture and hyperparameters. The embedding dimension is configured to 128, and the decoder comprises 6 attention layers. Each multi-head attention layer uses 8 heads, while the feed-forward layer dimension is set to 512. We configure the KNN to 100. All experiments are performed on a single NVIDIA GeForce RTX 3090 GPU with 24GB of memory.
@@
The evolutionary framework follows the approach proposed by EoH [liu2024evolution]. The population size is set to 20, and the evolution process is run for 105 generations. The parent heuristic is set to 2. The GPT-4o-mini model serves as the LLM backbone. For TSP/CVRP 1K, 5K, and 10K, we optimize one projection strategy for each scale, and generalize the strategy optimized on 10K to TSP/CVRP 50K and 100K due to the computational burden. At the beginning of the optimization, we input the projection strategy proposed in [fang2024invit] as a seed to guide the heuristic design.
@@
We compare TTPL with six categories of methods: Classical Solvers: Concorde [concorde], LKH3 [helsgaun2017extension] and HGS [vidal2022hybrid]; Insertion Heuristic: Random Insertion; Construction-based NCO Methods: POMO [kwon2020pomo], BQ [drakulic2023bq], LEHD [luo2023neural], INVIT [fang2024invit] and SIGD [pirnay2024self]; Heatmap-based Methods: DIFUSCO [sun2023difusco]; Decomposition-based Method: GLOP [ye2024glop], H-TSP [pan2023h]; Local Construction-based Method: ELG [gao2023towards].
@@
We evaluate the performance using the optimality gap and inference time. The optimality gap quantifies the percentage difference between solutions generated by our method and the ground truth obtained from LKH3 (TSP) or HGS (CVRP). Inference time, reported in seconds (s), minutes (m), and hours (h), captures the computational efficiency of generating solutions across varying problem scales. For TTPL, we report three types of results, those obtained by greedy trajectory with a purely designed projection method, those obtained by greedy trajectory with our designed projection and MVDF, and those obtained with projection, MVDF, and Random Re-Construction (RRC) under 1000 iterations [luo2023neural].